A) Solve the system of equations by substitution. B) Is this system consistent independent, consistent dependent or inconsistent? C)Describe the graph without drawing it
x-y=1
3x=3y+3

Answer by stanbon(75887)

A) Solve the system of equations by substitution.
x-y=1
3x=3y+3
-----
Modify for elimination::
y = x-1
y = x-1
------------------
Subtract to get 0 = 0
-------
The system is true for all points (x, x-1), where x is a Real Number.
------------

B) Is this system consistent independent, consistent, dependent or inconsistent?
Ans: dependent; consistent
================
C)Describe the graph without drawing it.
One line on top of the other.
---
Cheers,
Stan H.
======
--------